# Long-range angular correlations of particle displacements at a plastic-to-elastic transition in jammed amorphous solids

### v1

@@ Abstract
The data of $G^{*}$ and $\theta^{*}$ suffer strong finite-size effects and thus it is necessary to consider a proper finite-size scaling. We find that they can be collapsed when plotted as functions of a rescaled pressure, $\hat{p}=pN^{3/2}$ (Fig. 4c-d). To understand this finite-size scaling, recall that near the jamming transition, the isostatic length $\ell_{\rm iso}\sim(Z-Z_{\rm iso})^{-1}\sim p^{-1/3}$ sets a natural length scale, where $Z$ is the average coordination number and $Z_{\rm iso}=2d=4$ is the isostatic coordination number in $d=2$ dimensions [34]. According to the argument in Ref. [23], the PE transition occurs when $\ell_{\rm iso}$ matches the largest screening length $\ell_{\rm s}\sim 1/\kappa_{1}\approx r_{\rm out}/\pi\sim N^{1/2}$, which gives $p_{\rm c}\sim N^{-3/2}$, consistent with our data. Note that this scaling holds only for Hertzian systems. For Hookean systems, it becomes $p_{\rm c}\sim N^{-1}$ (Fig. S1).
Figure 5: Dependence of the PE transition on the inflation $d_{0}$. (a) $G^{*}$ and (b) $\theta^{*}$ as functions of $p/d_{0}^{3/4}$, for $N=67352$, where the dotted line marks $p_{\rm c}$. (c) The $d_{0}-p$ phase diagram (for $N=67352$), where the dotted line is $d_{0}\sim p_{\rm c}^{4/3}$.
Intuitively, the PE transition should also dependent on the amount of inflation, since a larger inflation could introduce stronger plasticity and non-linear effects. Examining the dependence of the PE transition on the inflation parameter $d_{0}$ reveals that $p_{\rm c}\sim d_{0}^{3/4}$. In Fig. 5a-b, we plot $\theta^{*}$ and $G^{*}$ obtained from inflation simulations with a few different $d_{0}$. The data can be collapsed as functions of $p/d_{0}^{3/4}$. To understand this scaling relationship, we assume that the effect of $d_{0}$ is controlled by the typical strain to create an additional contact by inflation $\epsilon^{\rm mk}$, i.e., $d_{0}\sim\epsilon^{\rm mk}$. Ref. [35] shows that the finite-size scaling of $\epsilon^{\rm mk}$ near jamming is, $\epsilon^{\rm mk}\sim 1/N^{2}$. Combining it with the finite-size scaling $p\sim N^{-3/2}$ discussed above, one obtains $d_{0}\sim N^{-2}\sim p^{4/3}$, consistent with the simulation data. Based on this analysis, a $d_{0}-p$ phase diagram is proposed (see Fig. 5c). The boundary between quasi-elastic and plastic regimes is $d_{0}\sim p_{\rm c}^{4/3}$. Interestingly, a similar strain-pressure phase diagram is conjectured in Ref. [36] (the two regimes are called linear and shear softening regimes there), although the linear regime could not be accessed in that experiment.
@@
Next we demonstrate that the system has long-range angular correlations at $\hat{p}_{\rm c}$, even though $\theta^{*}$ does not appear to diverge. Instead, in our setup, the long-range correlation corresponds to the convergence, $\theta^{*}\to\theta^{*}_{\rm c}\approx 0.3\pi$ and $G^{*}\to G^{*}_{\rm c}\approx 0.3$, when $\hat{p}\to\hat{p}_{\rm c}$ from below. To show that, we perform the Fourier transform of $d_{r}(\theta)$ and plot the (normalized) power spectrum $S(f)$ in Fig. 3b at different $\hat{p}$, for $N=67352$. Interestingly, $S(f)$ displays a power-law with an exponent $-1.5$ (which we term $1/f^{1.5}$ noise) at $\hat{p}_{\rm c}\approx 10^{8}$, indicating that the system is scale-free at the transition. The determined $\hat{p}_{\rm c}\approx 10^{8}$ is consistent with $\hat{p}_{\rm c}\approx 3\times 10^{8}$ estimated in Fig. 2c-d (considering strong data fluctuations there) and $\hat{p}_{\rm c}\approx 8\times 10^{7}$ reported in Ref. [23] (converted from $p_{\rm c}\approx 3.5$ for $N=80000$). In the analyses below we take $\hat{p}_{\rm c}\approx 10^{8}$ as the value of the rescaled transition pressure. In the Fig. S2, we show that the $1/f^{1.5}$ behavior at $\hat{p}_{\rm c}$ is robust against varying $\hat{r}$, and is universal in Hertzian and Hookean systems.
@@
where $f^{\dagger}$ is the only fitting parameter, and $A=1+(1/f^{\dagger})^{1.5}$ is fixed by $f^{\dagger}$ due to the normalization $S(f=1)=1$. Here $f^{\dagger}$ separates the white noise at $f<f^{\dagger}$ with $S(f)\sim 1$ from the $1/f^{1.5}$ noise at $f>f^{\dagger}$. Since $f^{\dagger}$ is a characteristic wave number, one expects a corresponding angle, $\theta^{\dagger}\sim 1/f^{\dagger}$. The physical meaning of $\theta^{\dagger}$ can be interpreted as follows: the angular correlation between displacements exists up to $\theta^{\dagger}$, beyond which the displacements are uncorrelated like white noise. Thus $\theta^{\dagger}$ is another measure of the correlation angle, and we expect $\theta^{\dagger}\sim\theta^{*}\sim 1/f^{\dagger}$ for large $f^{\dagger}$. Since one can simply assume that the displacements are uncorrelated in the fluid state, it is expected that $f^{\dagger}\to\infty$ (or $\theta^{\dagger}\to 0$) approaching the jamming transition, $\hat{p}\to\hat{p}_{\rm J}=0$. On the other hand, an elastic medium should have long-range correlated displacements, which means that $f^{\dagger}\to 0$ (or $\theta^{\dagger}\to\infty$), approaching the PE transition, $\hat{p}\to\hat{p}_{\rm c}$. These expectations are consistent with the data in Fig. 3b.
The divergence of $\theta^{\dagger}$, and its connection to $\theta^{*}$, can be established more rigorously. By performing the inverse Fourier transform of Eq. (5), it is easy to reconstruct numerically $G_{\rm GL}(\theta)$, from which $\theta_{\rm GL}^{*}$ can be determined by $G_{\rm GL}(\theta=\theta_{\rm GL}^{*})=0.1$ (see Fig. S3). This procedure gives a quantitative relation between $\theta^{*}_{\rm GL}$ and $f^{\dagger}$. For large $f^{\dagger}$, $\theta_{\rm GL}^{*}=k/f^{\dagger}$ with $k\approx 81$. According to this, we define $\theta^{\dagger}=k/f^{\dagger}$, which recovers the equivalence $\theta^{\dagger}=\theta^{*}$ at small $\hat{p}$ (Fig. 4d). Near $\hat{p}_{\rm c}$, $\theta^{\dagger}$ diverges such that,
@@
where $\mu\approx 0.9$ from fitting (see Fig. 4d-inset). On the other hand, $\theta^{*}$ approaches a finite value $\theta^{*}_{\rm c}=\theta^{*}_{\rm GL,c}\approx 0.3\pi$ near $\hat{p}_{\rm c}$. We conclude that the long-range angular correlation at $\hat{p}_{\rm c}$ is captured by the power-law divergence of $\theta^{\dagger}$ as in Eq. (6), while $\theta^{*}$ does not diverge due to the periodicity of $d_{r}(\theta)$ in $\theta\in[0,2\pi]$ (see Supplementary Information Sec. V for more discussions).
Discussion
This study opens several questions for future studies. In particular, the origin of the $f^{-1.5}$ spectrum at $p_{\rm c}$ needs to be recovered. To this end, it will be useful to generalize the current setup of circular inflation to other types of deformations [37, 38, 39]. Despite the similarity between the current problem and two-dimensional melting, unlike the latter, we expect that the two-step transitions discussed here would also appear in higher dimensions [21]. A heuristic argument shows that the hexatic phase transition in two-dimensional melting is a consequence of the competition between the vortex entropy $TS$ and energy $E$, both of which depend logarithmically on the system size in two dimensions. This argument can not be applied to the current granular system, since the temperature is zero. On the other hand, the match between the overlapping effects due to over-compression and the energy barriers required for plastic events (as discussed in the Introduction) may provide an alternative mechanism universal in any dimensions. Finally, it would be very interesting to reconcile the MFRT and MRST approaches.
Acknowledgments: We warmly thank Matteo Baggioli, Bulbul Chakraborty, Michael Moshe, Jin Shang, and Jie Zhang for inspiring discussions. We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 12047503 and 12404290), from Chinese Academy of Sciences (Grant ZDBS-LY-7017), and from Wenzhou Institute (Grant WIUCASQD2023009). IP acknowledges support from the ISF under grant #3492/21 (collaboration with China) and the Minerva Center for “Aging, from physical materials to human tissues” at the Weizmann Institute. In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.
@@ Discussion
I. Model
The two-dimensional model employed in our numerical simulations consists of $N$ equimolar frictionless bi-disperse disks with radii $R_{1}=0.5$ and $R_{2}=0.7$ (the length unit is the diameter of small disks). The normal force $\bm{\mathrm{F}}_{ij}$ between particles $i$ and $j$ is,
$\bm{\mathrm{F}}_{ij}^{(n)}=k_{n}\Delta_{ij}^{(n)}\hat{\bm{\mathrm{n}}}_{ij}-\gamma_{n}\bm{\mathrm{v}}_{ij}^{(n)},$ (S1)
where $\hat{\bm{\mathrm{n}}}_{ij}=\bm{\mathrm{r}}_{ij}/r_{ij}$ is the unit inter-particle distance, and $\bm{\mathrm{v}}_{ij}^{(n)}$ is the normal component of the relative velocity. The overlap between two contacting particles is defined as $\Delta_{ij}^{(n)}=R_{i}+R_{j}-\left|\bf{r}_{i}-\bf{r}_{j}\right|$. The spring coefficient is $k_{n}=k_{n}^{{}^{\prime}}\sqrt{\Delta_{ij}^{(n)}R_{ij}^{-1}}$ for the Hertzian interaction, and $k_{n}=k_{n}^{{}^{\prime}}$ for the Hookean interaction, with $k_{n}^{{}^{\prime}}=2\times 10^{5}$ and $R_{ij}^{-1}=R_{i}^{-1}+R_{j}^{-1}$. The damping coefficient is $\gamma_{n}=500$ for both models. The mass of all particles is set to be one. The presented data are for the Hertzian potential unless otherwise specified.
II. Simulation methods
The initially random configuration is generated under periodic boundary conditions at a fixed packing fraction $\varphi$. The system is then rapidly quenched to reach mechanical equilibrium, and the mechanical pressure $p$ is measured. After this step, simulations are performed in a circular area with a fixed boundary condition: a small particle is randomly chosen as the inflation center at $r=0$, and all particles at $r\geq r_{\rm out}$ are fixed.
The inflation simulation is carried out as follows. We first instantaneously inflate a small disk at the center by a factor of 1.2, i.e., $r_{\rm in}\to r_{\rm in}+d_{0}$ with $r_{\rm in}=0.5$, then minimize the energy of the entire system after the inflation. The inflation parameter $d_{0}=0.1$ unless otherwise specified. The position of the central particle is fixed during the minimization. For jammed configurations with $p>0$ and $\varphi>\varphi_{\rm J}$, the energy is dissipated by damped molecular dynamics (MD) simulations following the previous study [23]. The minimization procedure is terminated when the net force per particle $F_{\rm net}\leq 10^{-7}$. For unjammed configurations ($p=0$, $\varphi<\varphi_{\rm J}$), the damped MD simulation can create an inhomogeneous flow, which results in large voids that are empty of particles. To avoid such highly inhomogeneous displacements, the FIRE algorithm [40], which minimizes the energy more rapidly than the damped MD algorithm, is used for unjammed configurations. The above protocols are implemented using the LAMMPS package [41]. The presented data are averaged over more than 400 independent samples.
III. Finite-size scaling in Hookean systems
Let us generalize the argument given in the main text and consider the $p$ $N$ scaling for an inter-particle potential $U(\delta)\sim\delta^{\alpha}$ for an arbitrary $\alpha$ ($\delta$ is the dimensionless inter-particle overlap). The scaling $l_{\rm iso}\sim(Z-Z_{\rm iso})^{-1}$ is independent of $\alpha$ [34], while $(Z-Z_{\rm iso})\sim(\varphi-\varphi_{\rm J})^{1/2}\sim p^{\frac{1}{2(\alpha-1)}}$ depends on $\alpha$ [1, 32]. It follows that $p\sim p_{\rm c}\sim N^{-(\alpha-1)}$ for any $\alpha$. From this general scaling, we recover $p\sim N^{-3/2}$ for Hertzian systems ($\alpha=5/2$), and obtain $p\sim N^{-1}$ for Hookean systems ($\alpha=2$). The latter is confirmed by the simulation data presented in Fig. S1.
IV. Dependence of the power spectrum on the radius
@@
V. Inverse Fourier transform of the generalized Lorentzian spectrum
@@
where $B=\sum_{f}\frac{1}{1+(f/f^{\dagger})^{1.5}}$ is the normalizing factor, and $f$ is an integer because $d_{r}(\theta)$ is periodic in $\theta\in[0,2\pi]$. Following the main text, we define $\theta^{*}_{\rm GL}$ by $G_{\rm GL}(\theta=\theta^{*}_{\rm GL})=0.1$. As shown in Fig. S3, the curve $\theta^{*}_{\rm GL}(f^{\dagger})$ is close to the simulation data $\theta^{*}$ (for $\hat{p}<\hat{p}_{\rm c}$). In the large $f^{\dagger}$ limit (i.e., $\hat{p}\to 0$), $\theta^{*}_{\rm GL}(f^{\dagger}\to\infty)=k/f^{\dagger}$, with $k\approx 81$. Based on this, we define $\theta^{\dagger}=k/f^{\dagger}$ in the main text. In the small $f^{\dagger}$ limit (i.e., $\hat{p}\to\hat{p}_{\rm c}$), $\theta^{*}_{\rm GL}(f^{\dagger}\to 0)=\theta_{\rm GL,c}^{*}\approx 0.3\pi$. In addition, $G^{*}_{\rm GL,c}\equiv-G_{\rm GL}(\theta=\pi)\approx 0.3$.
@@
Figure S1: Finite-size scaling in Hookean disks. The simulation data of $G^{*}$ and $\theta^{*}$ can be collapsed as functions of (a,b) $pN$, but not (c,d) $pN^{3/2}$.

### v2

@@ Abstract
The data of $G^{*}$ and $\theta^{*}$ suffer strong finite-size effects and thus it is necessary to consider a proper finite-size scaling. We find that they can be collapsed when plotted as functions of a rescaled pressure, $\hat{p}=pN^{3/2}$ [Figs. 4(c) and  4(d)]. To understand this finite-size scaling, recall that near the jamming transition, the isostatic length $\ell_{\rm iso}\sim(Z-Z_{\rm iso})^{-1}\sim p^{-1/3}$ sets a natural length scale, where $Z$ is the average coordination number and $Z_{\rm iso}=2d=4$ is the isostatic coordination number in $d=2$ dimensions [39]. According to the argument in Ref. [23], the PE transition occurs when $\ell_{\rm iso}$ matches the largest screening length $\ell_{\rm s}\sim 1/\kappa_{1}\approx r_{\rm out}/\pi\sim N^{1/2}$, which gives $p_{\rm c}\sim N^{-3/2}$, consistent with our data (see SM  [28]). For a more general interaction potential $U(\delta)\sim\delta^{\alpha}$, $p_{\rm c}(N,d_{0})$ of different $N$ and inflation $d_{0}$ obeys a scaling function $N^{\beta}p_{\rm c}(N,d_{0})=\mathcal{G}(N^{2}d_{0})$, where $\beta=2(\alpha-1)$ and $\mathcal{G}(x)\sim x^{\beta/4}$ (see Appendix D and Sec. S5 in SM [28]).
@@
Next we demonstrate that the system has long-range angular correlations at $\hat{p}_{\rm c}$, even though $\theta^{*}$ does not appear to diverge. In our setup, the long-range correlation corresponds to the convergence, $\theta^{*}\to\theta^{*}_{\rm c}\approx 0.3\pi$ and $G^{*}\to G^{*}_{\rm c}\approx 0.3$, when $\hat{p}\to\hat{p}_{\rm c}$ from below. To show that, we perform the Fourier transform of $d_{r}(\theta)$ and plot the (normalized) power spectrum $S(f)$ in Fig. 3(b) at different $\hat{p}$, for $N=67352$. Interestingly, $S(f)$ displays a power-law with an exponent $-1.5$ (which we term $1/f^{1.5}$ noise) at $\hat{p}_{\rm c}\approx 10^{8}$, indicating angular scale-free behavior. The determined $\hat{p}_{\rm c}\approx 10^{8}$ is consistent with $\hat{p}_{\rm c}\approx 3\times 10^{8}$ estimated in Fig. 2 (considering strong data fluctuations there) and $\hat{p}_{\rm c}\approx 8\times 10^{7}$ reported in Ref. [23] (converted from $p_{\rm c}\approx 3.5$ for $N=80000$). In the analyses below we take $\hat{p}_{\rm c}\approx 10^{8}$ as the value of the rescaled transition pressure. In Sec. S6 of SM [28], we show that the $1/f^{1.5}$ behavior at $\hat{p}_{\rm c}$ is robust against varying $\hat{r}$, and is universal in Hertzian and Hookean systems.
@@
where $f^{\dagger}$ is the only fitting parameter, and $A=1+(1/f^{\dagger})^{1.5}$ is fixed by $f^{\dagger}$ due to the normalization $S(f=1)=1$. Here $f^{\dagger}$ separates the white noise at $f<f^{\dagger}$ with $S(f)\sim 1$ from the $1/f^{1.5}$ noise at $f>f^{\dagger}$. Since $f^{\dagger}$ is a characteristic wave number, one can define a corresponding angle, $\theta^{\dagger}\equiv k/f^{\dagger}$, with $k=0.45$ estimated in Sec. S7 of SM [28, 40]. The physical meaning of $\theta^{\dagger}$ can be interpreted as follows: the angular correlation between displacements exists up to $\theta^{\dagger}$, beyond which the displacements are uncorrelated like white noise. Assuming uncorrelated displacements in the fluid state, one expects that $\theta^{\dagger}\to 0$ approaching the jamming transition. On the other hand, the scale-free spectrum $S(f)$ at $\hat{p}_{\rm c}$ suggests that $\theta^{\dagger}\to\infty$ (i.e., $f^{\dagger}\to 0$) approaching the PE transition. Moreover, in the plastic phase, $S(f)$ deviates from the scale-free scaling due to a finite $\theta^{\dagger}$. These expectations are consistent with the data in Fig. 3(b). In particular, near $\hat{p}_{\rm c}$, $\theta^{\dagger}$ diverges such that
@@
where $\mu\approx 0.9$ from fitting [see Fig. 4(d) (inset)]. Eq. (6) is the angular counterpart of Eq. (3). Note that $\theta^{\dagger}$ diverges at $p_{\rm c}$ (or $\varphi_{\rm c}$) and $\xi$ diverges at $p_{\rm J}$ (or $\varphi_{\rm J}$).
The two correlation angles, $\theta^{*}$ and $\theta^{\dagger}$, coincide at small $\hat{p}$ [Fig. 4(d)]. The angle $\theta^{*}$ does not diverge at $p_{\rm c}$, because it is defined in a finite domain $\theta\in[0,\pi]$. On the other hand, $\theta^{\dagger}$ can diverge due to the analytic extension of the frequency, $f^{\dagger}=k/\theta^{\dagger}\to 0$. In linear inflation geometry, we expect that $S(f)\sim f^{-1.5}$ would still hold at $p_{\rm c}$ and the associated length scale perpendicular to the inflation direction would diverge without the need of analytic extension.
Discussion. The angular behavior of the particle displacements $d_{r}(\theta)$ in response to the central inflation can be summarized as follows. (i) In the unjammed phase ($p=0$), the displacements are uncorrelated and white-noiselike, with a zero correlation angle. (ii) In the plastic phase ($0<p<p_{\rm c}$), the correlation angle is finite, and the displacements are correlated only within the correlation angle. (iii) In the quasielastic phase ($p>p_{\rm c}$), the correlation angle is finite because the displacement field is dominated by dipolar and quadrupolar patterns. The long-range angular correlation emerges at the PE transition ($p_{\rm c}$), captured by a diverging correlation angle $\theta^{\dagger}$ and a power-law spectrum $S(f)\sim f^{-1.5}$.
This study opens several questions for future studies. In particular, the origin of the $f^{-1.5}$ spectrum at $p_{\rm c}$ needs to be recovered. It will be useful to generalize the current setup of circular inflation to other types of deformations [41, 42, 43]. Despite the similarity between the current problem and two-dimensional melting, unlike the latter, we expect that the two-step transitions discussed here would also appear in higher dimensions [21]. Finally, it would be very interesting to reconcile the MFRT and MRST approaches.
Acknowledgments. We warmly thank Matteo Baggioli, Bulbul Chakraborty, Wouter Ellenbroek, Wenyue Fan, Michael Moshe, Jin Shang, and Jie Zhang for inspiring discussions. We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 12047503 and 12404290), from Chinese Academy of Sciences (Grant ZDBS-LY-7017), and from Wenzhou Institute (Grant WIUCASQD2023009). IP acknowledges support from the ISF under grant #3492/21 (collaboration with China) and the Minerva Center for “Aging, from physical materials to human tissues” at the Weizmann Institute. In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.
@@
S1 Dipolar and quadrupolar patterns of the displacement fields at $p>p_{\rm c}$ under square periodic boundary conditions
In the quasi-elastic regime ($p>p_{\rm c}$), dipolar and quadrupolar patterns of the displacement fields are commonly observed (see Fig. 1(g,j)). To examine if such anisotropic patterns are caused by the fixed circular boundary conditions, we perform additional inflation simulations in a square simulation box with periodic boundary conditions (the center of mass of the entire system is fixed). As shown in Fig. S1, dipolar and quadrupolar patterns are repeatedly observed in different samples, confirming that the anisotropic patterns at $p>p_{\rm c}$ do not significantly depend on the boundary conditions.
Figure S1: Radial displacement fields under periodic boundary conditions. Presented in (a-j) are the heap maps of $\hat{d}_{r}(x,y)$ obtained for ten independent samples of $N=2774$ disks at $p=3.5\times 10^{3}$.
S2 Ensemble-averaged radial displacements
The data of rescaled radial displacements $\hat{d}_{r}(r)$ presented in Fig. 1(b,e,h,k) are obtained for single configurations. The single-sample curves are suitable to fit the solutions of the mechanical response screening theory (MRST). Previous inflation simulations have also considered ensemble-averaged quantities, which are averaged over many independent configuration samples [29, 30]. The ensemble-averaged $\overline{\hat{d}_{r}(r)}$ are presented in Fig. S2, for a few different pressures $p$ and inflation parameters $d_{0}$. For the largest $p$ and smallest $d_{0}$, the behavior of $\overline{\hat{d}_{r}(r)}$ is close to the elastic solution Eq. (1). With decreasing $p$ or increasing $d_{0}$, deviations from Eq. (1) are observed. When $p$ is near or below $p_{\rm c}$, or when $d_{0}$ is near or larger than $d_{\rm c}$, the ensemble-averaged $\overline{\hat{d}_{r}(r)}$ mix samples with elastic and anomalous responses, and thus they can be fitted by neither Eq. (1) nor Eq. (3). In order to compare with theoretical solutions (Eq. (1) or (3)), one should use single-sample data $\hat{d}_{r}(r)$ instead of the ensemble-averaged data $\overline{\hat{d}_{r}(r)}$. Fig. S2 also shows that the coarse-grained elasticity  [29, 30] works in the quasi-elastic phase ($p>p_{\rm c}$ and $d_{0}<d_{\rm c}$), but fails in the plastic phase ($p<p_{\rm c}$ and $d_{0}>d_{\rm c}$) where nonlinear behavior is important.
Figure S2: Ensemble-averaged $\overline{\hat{d}_{r}(r)}$ for $N=67352$. (a) A few different pressures $p$, for a fixed $d_{0}=0.1$. (b) A few different $d_{0}$, for a fixed $p=9\times 10^{-1}$.
S3 Jamming transition
The unjammed configurations with $p=0$ are studied and it is found that the displacement response of the central inflation propagates only up to a finite distance. Without loss of generality, we define the length scale $\xi$ by $d_{r}(r=\xi)=d_{\rm th}=10^{-3}$. The power-law divergence Eq. (3) of $\xi$ with $\nu\approx 0.8$ is confirmed in Fig. S3 when approaching $\varphi_{\rm J}=0.842$, consistent with previous studies  [38, 34]. The exponent $\nu$ is unchanged when the threshold $d_{\rm th}$ in the definition of $\xi$ is decreased from $10^{-3}$ to $10^{-6}$ (see Fig. S4).
Figure S3: Jamming transition. The characteristic length $\xi$ for the jamming transition as a function of (a) the packing density $\varphi$, and (b) $\varphi_{\rm J}-\varphi$, for a few different $N$.
Figure S4: Correlation length $\xi$ obtained using different thresholds. The threshold $d_{\rm th}$ in the definition $d_{r}(r=\xi)=d_{\rm th}$ is respectively (a) $10^{-4}$, (b) $10^{-5}$ and (c) $10^{-6}$. The exponent $\nu\approx 0.8$ is unchanged.
S4 Dependence of $\theta^{*}$ on the threshold $G_{\rm th}$
Figure S5 shows that the finite-size scaling $p_{\rm c}\sim N^{-3/2}$ in Hertzian systems is nearly independent of the threshold value $G_{\rm th}$ used in the definition of $\theta^{\ast}$, $G(\theta=\theta^{\ast})=G_{\rm th}$.
Figure S5: Correlation angle $\theta^{*}$ as a function of $pN^{3/2}$, using different threshold values $G_{\rm th}$. In (a-c), $G_{\rm th}=0.02,0.05,0.2$ respectively.
S5 Additional data on the scaling analysis in Hookean systems
In Hertzian systems, we have shown the $N$-dependence of $p_{\rm c}$, $p_{\rm c}\sim N^{-3/2}$, in the main text. For Hookean disks, the numerical results suggest a finite-size scaling $p\sim N^{-1}$, instead of $p\sim N^{-3/2}$, for a fixed $d_{0}$ (see Fig. S6). In addition, the data of $\theta^{*}$ and $G^{*}$ can be collapsed as functions of $pN/d_{0}^{1/2}$ (see Fig. S7).
Figure S6: Finite-size scaling in Hookean disks. The simulation data of $G^{*}$ and $\theta^{*}$ can be collapsed as functions of (a,b) $pN$, but not (c,d) $pN^{3/2}$. The inflation $d_{0}=0.1$ is fixed.
Figure S7: Dependence of the PE transition on the inflation $d_{0}$ and the system size $N$, for Hookean systems. (a) $G^{*}$ and (b) $\theta^{*}$ as functions of $pN/d_{0}^{1/2}$, for three different $N$ and four different $d_{0}$, where the dotted line marks $p_{\rm c}N/d_{0}^{1/2}$. (c) The $N$-rescaled $d_{0}-p$ phase diagram, where the dotted line is $(N^{2}d_{0})\sim(N^{2}p_{\rm c})^{2}$.
S6 Dependence of the power spectrum on the radius
@@
S7 Inverse Fourier transform of the generalized Lorentzian spectrum
@@
where $B=\sum_{f}\frac{1}{1+(f/f^{\dagger})^{1.5}}$ is the normalizing factor, and $f$ is an integer because $d_{r}(\theta)$ is periodic in $\theta\in[0,2\pi]$. Following the main text, we define $\theta^{*}_{\rm GL}$ by $G_{\rm GL}(\theta=\theta^{*}_{\rm GL})=0.1$. As shown in Fig. S9, the curve $\theta^{*}_{\rm GL}(f^{\dagger})$ is close to the simulation data $\theta^{*}$ (for $\hat{p}<\hat{p}_{\rm c}$). In the large $f^{\dagger}$ limit (i.e., $\hat{p}\to 0$), $\theta^{*}_{\rm GL}(f^{\dagger}\to\infty)=k/f^{\dagger}$, with $k\approx 0.45$. Based on this, we define $\theta^{\dagger}=k/f^{\dagger}$ in the main text. In the small $f^{\dagger}$ limit (i.e., $\hat{p}\to\hat{p}_{\rm c}$), $\theta^{*}_{\rm GL}(f^{\dagger}\to 0)=\theta_{\rm GL,c}^{*}\approx 0.3\pi$. In addition, $G^{*}_{\rm GL,c}\equiv-G_{\rm GL}(\theta=\pi)\approx 0.3$.
@@
S8 Isostatic length
The isostatic length $\ell_{\rm iso}$ has been determined explicitly in previous central inflation simulations, using the fluctuation data of the change in contact force, $\Delta F$ [29, 30]. For a pair of contacting particles $i$ and $j$, by definition the force change $\Delta F_{ij}$ is proportional to $d_{\parallel,ij}$, i.e., $\Delta F_{ij}\equiv-k_{n}d_{\parallel,ij}$, where $d_{\parallel,ij}$ is the parallel component of the relative displacement $\bm{\mathrm{d}}_{ij}=\bm{\mathrm{d}}_{j}-\bm{\mathrm{d}}_{i}$ along the contact bond, and $k_{n}$ is the spring coefficient. Inspired by this method, here we use the fluctuations of $d_{\parallel}$ to extract $\ell_{\rm iso}$. Note the difference between $d_{r}$ and $d_{\parallel}$: $d_{r}$ is the projection of the displacement $\bm{\mathrm{d}}$ onto the radial direction with respect to the inflation center of the simulation system; $d_{\parallel}$ is the projection of the relative displacement $\bm{\mathrm{d}}_{ij}$ onto the bond direction, which represents the change in the bond length. Comparing the visualized fields of $d_{\parallel}(r,\theta)$ in Fig. S10 with the corresponding $d_{r}(r,\theta)$ in Fig. 1(d, g, j), one sees that $d_{\parallel}(r,\theta)$ captures the physics of isostatic stability, rather than that of the plastic screening effects. The $d_{\parallel}(r,\theta)$ field looks disordered within a length scale $\ell_{\rm iso}$ from the inflation center. At larger length scales beyond $\ell_{\rm iso}$, the role of disorder is smeared out and the $d_{\parallel}(r,\theta)$ field looks more similar to what is expected for a continuous elastic medium. In other words, $d_{r}(r,\theta)$ and $d_{\parallel}(r,\theta)$ encode respectively the information of the screening length $\ell_{\rm s}$ and the isostatic length $\ell_{\rm iso}$.
In order to evaluate $\ell_{\rm iso}$ quantitatively, we follow Refs. [29, 30] and compute the root mean square fluctuations $h_{\parallel}(r)=\sqrt{\left\langle\left(d_{\parallel}(r)-\langle d_{\parallel}(r)\rangle\right)^{2}\right\rangle}$, where the average $\langle x\rangle$ is taken over along the circle of radius $r$. As shown in Fig. S11, the data of $h_{\parallel}(r)$ at different $p$ all collapse when plotted as functions of $r\Delta Z$ with $\Delta Z=Z-Z_{\rm iso}$, suggesting the characteristic length scale $\ell_{\rm iso}\sim 1/\Delta Z$. The tail of $h_{\parallel}(r)$ follows a power-law $h_{\parallel}(r)\sim r^{-1}$, corresponding to the regime where the effects of disorder disappear. The threshold to reach this asymptotic power-law defines a proper pre-factor for the scaling of $\ell_{\rm iso}$ (see Fig. S11b),
$\ell_{\rm iso}=\frac{C}{\Delta Z},$ (S2)
with $C\approx 6$, consistent with Refs. [29, 30]. We have checked that the fluctuations of $d_{r}(r)$ do not collapse as functions of $r\Delta Z$ (see Fig. S11(c,d)).
Figure S10: Visualization of the $d_{\parallel}$ field at (a) $p=4.4\times 10^{-2}$, (b) $p=5.2\times 10^{2}$ and (c) $p=8.6\times 10^{2}$, for $N=67452$ and $d_{0}=0.1$. The red (blue) lines indicate compressed (stretched) bonds with $d_{\parallel}<0$ ($d_{\parallel}>0$). The black circles indicate $\ell_{\rm iso}$.
Figure S11: Fluctuations of $d_{\parallel}$. (a) The root mean square fluctuations $h_{\parallel}(r)$ as functions of $r\Delta Z$ for different $p$. The deviations from the master curve at large $r$ are due to boundary effects. (b) $h_{\parallel}(r)\times(r\Delta Z)$ as functions of $r\Delta Z$, showing a plateau at large distances. We choose a threshold $h_{\parallel}(r)\times(r\Delta Z)=1.3$ (dashed line) that gives the crossover value $\ell_{\rm iso}\Delta Z=C\approx 6$. For comparison, the fluctuations of $d_{r}(r)$, defined by $h_{r}(r)=\sqrt{\left\langle\left(d_{r}(r)-\langle d_{r}(r)\rangle\right)^{2}\right\rangle}$ are plotted in (c,d), which do not collapse as functions of $r\Delta Z$.
S9 Contact changes
Figure S12 shows that the inflation of the central particle causes multiple contact changes in our simulations. Thus the plastic-to-elastic (PE) transition does not correspond to the first contact change event analyzed in previous studies [24, 25].
Figure S12: Contact changes due to (a-c) inflation and (d-f) deflation of the central disk. The pressures are (a,d) $p=4.4\times 10^{-2}$, (b,e) $p=5.2\times 10^{2}$, (c,f) $p=8.6\times 10^{2}$ (corresponding to the second to fourth rows in Fig. 1). The inflation-deflation simulations are performed following the procedure $d_{0}=0\to 0.1\to 0$, and the configurations are equilibrated after the inflation and deflation. The contact networks before and after the inflation/deflation are compared. The broken and created contacts are marked by blue and red points respectively. Note that there is no contact change for unjammed systems since $Z=0$.